OpenStudy (ofmiceparade):

For the function f(x) = –2(x + 3)2 − 1, identify the vertex, domain, and range. The vertex is (3, –1), the domain is all real numbers, and the range is y ≥ –1. The vertex is (3, –1), the domain is all real numbers, and the range is y ≤ –1. The vertex is (–3, –1), the domain is all real numbers, and the range is y ≤ –1. The vertex is (–3, –1), the domain is all real numbers, and the range is y ≥ –1.

OpenStudy (ofmiceparade):

@Vocaloid por favor

Vocaloid (vocaloid):

given the function y = (x-h)^2 + k the vertex is (h,k)

OpenStudy (ofmiceparade):

.... so its A?

Vocaloid (vocaloid):

read carefully if (x-h) = (x+3), what is h?

OpenStudy (ofmiceparade):

3

Vocaloid (vocaloid):

nope, think carefully x - h = x + 3 subtract x from each side, what do you get?

OpenStudy (ofmiceparade):

-3?

OpenStudy (ofmiceparade):

You divide it by negative h?

Vocaloid (vocaloid):

right, so h = -3 and our vertex is (-3,-1)

Vocaloid (vocaloid):

now, to find the range, we can see that the graph is a downward parabola, so our range is y <= -1

OpenStudy (ofmiceparade):

okay

OpenStudy (ofmiceparade):

wouldn't it be C?

Vocaloid (vocaloid):

good, yes

OpenStudy (ofmiceparade):

omg thank you Senpai

OpenStudy (ofmiceparade):

I gave you medal for being Kawaii

Vocaloid (vocaloid):

pfff lol

OpenStudy (ofmiceparade):

Much fan

Vocaloid (vocaloid):

thanks ^_^